OpenStudy (anonymous):

Do the equations x = 4y + 1 and x = 4y – 1 have the same solution? How might you explain your answer to someone who has not learned algebra?

jimthompson5910 (jim_thompson5910):

x = 4y + 1 x - 4y = 1 --------------- x = 4y – 1 x - 4y = -1

jimthompson5910 (jim_thompson5910):

So x = 4y + 1 and x = 4y – 1 turns into x - 4y = 1 and x - 4y = -1